OpenStudy (anonymous):

What would make the de Broglie wavelength of some particle gbe large enough to be observed in an experiment? Large/ small velocity, or large/ small mass?

OpenStudy (preetha):

Sorry, got logged out of the internet. So the wavelength is inversely proportional to mass. Large objects have negligible wavelength. THat is why you and me dont have a wavelenght. And small objects like electrons have wavelengths.

OpenStudy (anonymous):

shouldnt it be like: wavelength \(\propto1/\)momentum momenum includes both mass and velocity

OpenStudy (preetha):

Yes. λ = hv/mv2 = h/mv lambda = wavelenght, mv2 = momentum. See: http://chemwiki.ucdavis.edu/Physical_Chemistry/Quantum_Mechanics/De_Broglie_Wavelength

OpenStudy (anonymous):

right. so, both gotta change

OpenStudy (anonymous):

or more precisely, both mass and velo. "can" be changed